Restart UT 09-09-15

OSA: Geno Bechetti

ISp: Elliott Shield

Telescope Manager: Jeff Urban

ISAs: Olga Kuhn, Barry Rothberg, J. Hill, D. Thompson (all from Tucson)

UT 02:20 - Socked in with thick clouds and rain. D. Thompson running LUCI-2 MOS checkout.

-Serious issues with the chillers, LUCI-2 being monitored. Repair work will be needed in the morning

-- %USERSIG{BarryRothberg - 2015-09-09}%

Comments

 
Topic revision: r2 - 09 Sep 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback